a transformer is used to convert 1200V AC into 48V AC by having a secondary coil that has 35 turns. The resistance of the primary coil is 5 ohms

a) What type of transformer is this?

b) What is the current in the primary coils?

c) How many turns are in the primary?

d) What is the power tht is running in the transformer?

e) What is the current in the secondary coil?
Answer #1

a) Step Down Transformer
b) V = IR
1200 = I * 5
I = 240A
c) V1/V2 = N1/N2
   N1 = N2 * V1 / V2
   = 35 * 1200 / 48
   = 875 Turns

d ) Power = V1 * I1 = V2 * I2
= 1200*240 W
   = 288000 W
e) I2 = V1 * I1 / V2
   = 6000 A